Let q(x1, x2)= 3x12 - 6x1x2 +11x22. Then find an orthonormal change of coordinate that diagonalizes the above quadratic form q


1
Expert's answer
2021-01-12T15:09:12-0500

 q(x1, x2)= 3x12 - 6x1x2 +11x22

matrix of the quadratic equation is "A=\begin{pmatrix}
 3 & -3 \\
 -3 & 11
\end{pmatrix}"


first we find the orthonormal diagonalization

"det(A -\lambda\Iota) = 0"


"\Iota = \begin{bmatrix}
 1 & 0 \\
 0 & 1
\end{bmatrix}"


"\lambda\Iota = 
\begin{bmatrix}
 \lambda\ & 0 \\
 0 & \lambda
\end{bmatrix}"

"\because A -\lambda\Iota = A=\begin{bmatrix}
 3 & -3 \\
 -3 & 11
\end{bmatrix} - \begin{bmatrix}
 \lambda\ & 0 \\
 0 & \lambda
\end{bmatrix} = \begin{bmatrix}
 3-\lambda\ & -3 \\
 -3 & 11-\lambda
\end{bmatrix}"


"det(A -\lambda\Iota) = [(3-\lambda)(11-\lambda)] - 9 =0"


"33 -3\lambda-11\lambda + \lambda^2 -9= 0"

"\lambda^2 - 14\lambda +24= 0"

"\lambda = 2 , \lambda =12" (eigen values)

when "\lambda=2: N(A-(2)\Iota)"

"\begin{bmatrix}
 1 & -3 \\
 -3 & 9
\end{bmatrix} \to" "\begin{bmatrix}
 1 & -3 \\
 0 & 0
\end{bmatrix}" eigen space : span "\begin{bmatrix}
 3 \\
 1 
\end{bmatrix}"


we normalize the vectors to get orthonormal vectors "{1 \over \sqrt{1 + 9}} \begin{bmatrix}
 3 \\
 1 
\end{bmatrix}"

"\begin{bmatrix}
 {3 \over \sqrt{10 }} \\
 {1 \over \sqrt{10}}
\end{bmatrix}"


when "\lambda=12: N(A-(12)\Iota)"

"\begin{bmatrix}
 -9 & -3 \\
 -3 & -1
\end{bmatrix} \to" "\begin{bmatrix}
 -9 & -3 \\
 0 & 0
\end{bmatrix}" eigen space : span"\begin{bmatrix}
 1 \\
 -3 
\end{bmatrix}"

we normalize the vectors to get orthonormal vectors "{1 \over \sqrt{1 + 9}} \begin{bmatrix}
 1 \\
 -3 
\end{bmatrix}"

"\begin{bmatrix}
 {1 \over \sqrt{10 }} \\
 {-3 \over \sqrt{10}}
\end{bmatrix}"


orthonormal diagonalization:

D= "\begin{bmatrix}
 2 & 0 \\
 0 & 12
\end{bmatrix},P= \begin{bmatrix}
 {3 \over \sqrt{10 }} & {1 \over \sqrt{10 }} \\
 {1 \over \sqrt{10 }} & -{3 \over \sqrt{10 }}
\end{bmatrix}"


Now we make a change of coordinates: Before it was q(y1,y2)= P-1x = PTx Now it will be q(x1,x2)= PT(y1,y2)

(x1,x2) = "\begin{bmatrix}
 {3 \over \sqrt{10 }} & {1 \over \sqrt{10 }} \\
 {1 \over \sqrt{10 }} & -{3 \over \sqrt{10 }}
\end{bmatrix}" (y1,y2)